OpenStudy (anonymous):

3log2(x-1)+log2(4)=5

OpenStudy (anonymous):

log2(4)=log2(2^2)=2 3log2(x-1)+2=5 3log2(x-1)=3 log2(x-1)=1 generally logx(x)=1 therefore log2(2)=1 x-1=2 x=3

OpenStudy (anonymous):

3log2 (x-1)+log2 4=5 log2[4*(x-1)^3] = 5 4(x-1)^3 = 2^5 = 32 (x-1)^3 = 8 x-1 = 2 x = 3

OpenStudy (tkhunny):

Have you considered logarithm rules? \(a\cdot log(b) = log\left(b^{a}\right)\) \(log(a) + log(b) = log(ab)\) \(log_{b}(a) = c \iff b^{c} = a\)

OpenStudy (anonymous):

GirlByte - how did you get the 32 in your answer?